OpenStudy (anonymous):

The table shows the height of a plant as it grows. What equation in point-slope form gives the plant’s height at any time? Let y stand for the height of the plant in cm and let x stand for the time in months.

OpenStudy (anonymous):

A. y – 15 = (x – 3) B. y – 15 = 5(x – 3) C. y – 3 = (x – 15) D. The relationship cannot be modeled.

OpenStudy (goformit100):

C. y – 3 = (x – 15)